Martina leaves home at 9am bicycling at a rate of 24 miles per hour, 2 hours later John leaves driving at the rate of 48 miles per hour. At what time will john catch up with Martina?
Answer by ankor@dixie-net.com(22740)

Martina leaves home at 9 am bicycling at a rate of 24 miles per hour, 2 hours later John leaves driving at the rate of 48 miles per hour. At what time will john catch up with Martina?
:
Let t = M's travel time
Then
(t-2) = J's travel time
:
One thing to remember on all these "catch-up" problems is that they will have traveled the same distance when that occurs. Write a distance equation from this fact.
:
Distance = speed * time
:
J's dist = M's dist
48(t-2) = 24t
48t - 96 = 24t
48t - 24t = +96
24t = 96
t = 96/24
t = 4 hrs is M's travel time
:
Find the time: 9:00 AM = 4:00 = 1 PM is when J catches up with M
:
Confirm our solution by finding the distances:
24*4 = 96 mi
48*(4-2)) = 96 mi
;
Did this make sense? Any questions?